# Scaling limit of heavy-tailed nearly unstable cumulative INAR($\infty$) processes and rough fractional diffusions

### v6

@@ Abstract
Scaling limit of heavy-tailed nearly unstable cumulative INAR($\infty$) processes and rough fractional diffusions
Yingli Wang  [^1] School of Mathematics, Shanghai University of Finance and Economics, Shanghai, People’s Republic of China; 2022310119@163.sufe.edu.cn., Chunhao Cai  [^2] School of Mathematics (Zhuhai), Sun Yat-Sen University, Zhuhai, People’s Republic of China; caichh9@mail.sysu.edu.cn, Ping He  [^3] School of Mathematics, Shanghai University of Finance and Economics, Shanghai, People’s Republic of China; pinghe@mail.shufe.edu.cn, Qinghua Wang  [^4] School of Mathematics, Shanghai University of Finance and Economics, Shanghai, People’s Republic of China; wangqinghua@mail.shufe.edu.cn
Keywords: heavy tailed unstable INAR($\infty$) processes, fractional CIR processes, rough volatility, option pricing.
MSC Classification: 60G22, 60F05
@@ 1 Introduction
Let $(\epsilon_{n})_{n\in\mathbb{Z}}$ be i.i.d. Poisson($\nu$) random variables. Let $(\eta_{k})_{k\geq 1}$ be a sequence of positive numbers in $\ell^{1}$. For $n\in\mathbb{Z}$, $k\geq 1$, and $l\geq 1$, let $\xi_{l}^{(n,k)}$ be mutually independent random variables with a Poisson($\eta_{k}$) distribution, which are also independent of the sequence $(\epsilon_{n})_{n\in\mathbb{Z}}$.
@@
To establish a more compact notation, we introduce the reproduction operator, denoted by $\circ$”. For a reproduction parameter $\alpha\geq 0$, we define the offspring sequence $(\xi_{n}^{(\alpha)})_{n\geq 1}$ as a sequence of i.i.d. random variables, where each element $\xi_{n}^{(\alpha)}$ is an offspring variable following a Poisson distribution with mean $\alpha$.
@@
$\alpha\circ Y:=\sum_{n=1}^{Y}\xi_{n}^{(\alpha)}$
provided that $Y$ is independent of the offspring sequence $(\xi_{n}^{(\alpha)})$. Using this operator, the system of stochastic difference equations can be compactly rewritten as:
$\epsilon_{n}=X_{n}-\sum_{k=1}^{\infty}\eta_{k}\circ X_{n-k}.$
Finally, we define $\nu$ as the immigration parameter, $(\epsilon_{n})$ as the immigration sequence, and $\eta_{k}$ as the reproduction coefficient for each non-negative integer $k$.
@@
$N_{n}=\sum_{s=1}^{n}X_{s}.$ (1)
Hawkes processes, introduced by Hawkes (1971), are continuous-time self-exciting point processes widely used in various fields. Discrete-time analogs, such as cumulative INAR($\infty$) processes, provide similar modeling capabilities, particularly for count data observed at fixed time intervals. However, the scaling limits of heavy-tailed, unstable cumulative INAR($\infty$) processes differ considerably from those of continuous-time counterparts, which warrants a dedicated investigation. Under certain conditions, the Poisson autoregressive process can be viewed as an INAR($\infty$) process with Poisson offspring. For a comprehensive discussion on Poisson autoregressive models and their connections to INAR and Hawkes processes, refer to Fokianos (2024), Huang and Khabou (2023) and Kirchner (2016). From Proposition 3 in Kirchner (2016), we see that if an INAR($\infty$) process $(X_{n})_{n\geq 1}$ starts from time $1$, it can also be defined by:
$\lambda_{n}:=\nu+\sum_{s=1}^{n-1}\eta_{n-s}X_{s},$ (2)
where $\nu>0$ is the immigration rate, and $(\eta_{n})_{n\geq 1}\in\ell^{1}$ represents the offspring distribution with $\eta_{n}\geq 0$ for all $n\in\mathbb{N}$. Given the history $\mathcal{F}_{n-1}$, the count $X_{n}$ follows a Poisson distribution with parameter $\lambda_{n}$, i.e.,
$X_{n}\mid\mathcal{F}_{n-1}\sim\text{Poisson}(\lambda_{n}).$
@@
under the skorohod topology, with $\sigma$ an explicit constant and $(W_{t})$ a Brownian motion. This result has been extended in Zhu (2013) to the case of nonlinear Hawkes processes. Regarding unstable Hawkes processes, the light-tailed case and the heavy-tailed case are respectively studied in Jaisson and Rosenbaum (2015) and Jaisson and Rosenbaum (2016). The light-tailed case corresponds to a CIR limit obtained under the crucial assumption
@@
a sequence of appropriately scaled random step functions formed from the process converges weakly towards a squared Bessel process. This result can be extended to INAR($\infty$) processes, which corresponds to what we term the “light-tailed” case. Consequently, the heavy-tailed case has not been addressed in the literature, a gap this paper aims to fill.
In this paper, we aim to provide a discrete-time counterpart to the continuous-time results found in Jaisson and Rosenbaum (2016), focusing on heavy-tailed unstable INAR($\infty$) processes. Specifically, we investigate an INAR($\infty$) process with
@@
where the constant $\alpha\in\left(\frac{1}{2},1\right)$. Inspired by the continuous-time case in Jaisson and Rosenbaum (2016), we consider a sequence of INAR processes $(X_{n}^{T})_{n\in\mathbb{N}}$ and analyze the asymptotic behavior as $T\to\infty$. The sequence of parameters $(\lambda_{n}^{T})$ is defined by
$\lambda_{n}^{T}=\mu^{T}+\sum_{s=1}^{n-1}\eta_{n-s}^{T}X_{s}^{T},\quad n\in\mathbb{N},$ (3)
where $\mu^{T}$ is a sequence of positive real numbers. As in Jaisson and Rosenbaum (2016), we establish specific assumptions on $\eta^{T}$ to facilitate our analysis. We establish the tightness of the renormalized cumulative INAR($\infty$) process $Y^{T}$ and associated martingale $Z^{T}$ defined in (5) and (7) which will be given below under specific assumptions, and show that any limit point $(Z,Y)$ of $(Z^{T},Y^{T})$ converges to a rough fractional diffusion process characterized by an integrated fractional Cox-Ingersoll-Ross (CIR) process.
@@
    A Computationally Superior Simulation Framework: Perhaps the most significant practical contribution is that INAR($\infty$) processes are substantially easier and more efficient to simulate than their continuous-time, singular-kernel Hawkes counterparts. Simulating our model only requires iterative sampling from a Poisson distribution. Our scaling limit theorem thus serves as a crucial bridge: it rigorously validates that this simple, efficient discrete simulation converges to the complex continuous process of interest. This directly addresses the question of “why not just use a Hawkes process?”: The INAR($\infty$) process provides a computationally tractable pathway for its analysis and application.
@@
    New Applications in Financial Modeling: This framework opens the door to new applications not explored in the original Hawkes process literature. For instance, as demonstrated in the recent work of Wang and Cui (2025), a bivariate version of this INAR($\infty$) framework can be shown to converge to the rough Heston model. This provides not only a novel microstructural foundation for rough volatility but also a highly efficient simulation engine for pricing both European and path-dependent options, a task that remains challenging for other methods.
Our primary result, encapsulated in Theorem 1, demonstrates that under heavy-tailed and unstable conditions, the cumulative INAR($\infty$) process converges to a rough fractional diffusion process. This finding is achieved by leveraging discrete renewal theory and martingale central limit theorems.
@@ 2 Assumptions and Main Results
First, we define an unstable cumulative INAR process with $\|\eta\|_{1}=1$. We propose two assumptions similar to those in Jaisson and Rosenbaum (2016).
Assumption 1.
@@
    The baseline offspring distribution $(\eta_{n})_{n\geq 1}$ has a heavy tail, characterized by the asymptotic behavior:
    $\eta_{n}\sim\frac{K}{n^{1+\alpha}}\quad\text{as}\quad n\to\infty,$
    for some $\alpha\in(\frac{1}{2},1)$ and some positive constant $K$.
@@
The assumption on the asymptotic behavior of $\eta_{n}$ in Assumption 1 is a standard condition for regularly varying sequences. A direct consequence of this assumption, based on Karamata’s Tauberian Theorem for power series, is that the tail sum of the sequence also has a well-defined asymptotic form. Specifically, it implies that:
@@
This equivalent formulation of the tail behavior is what we primarily use in our subsequent proofs.
Assumption 2.
@@
where $\delta=K\frac{\Gamma(1-\alpha)}{\alpha}$. The sequence $(\mu^{T})$ is shown in (3).
We first introduce some notation. Define
$M_{n}^{T}:=\sum_{s=1}^{n}\left(X_{s}^{T}-\lambda_{s}^{T}\right)=N_{n}^{T}-\sum_{s=1}^{n}\lambda_{s}^{T}.$
We also set $(A_{n}^{T})_{n\geq 1}\in\ell^{1}$ as the sequence defined on $\mathbb{N}$ by
$A_{n}^{T}:=\sum_{k=1}^{\infty}(\eta^{T})_{n}^{*k},$ (4)
where * denotes the discrete convolution, which, for two non-negative sequences $(q_{n})_{n\geq 1}$, $(m_{n})_{n\geq 1}\in\ell^{1}$,
@@
and $\eta^{*k+1}$ denotes the discrete convolution of $\eta^{*k}$ with $\eta$, i.e. $\eta^{*k+1}=\eta*\eta^{*k}$. $(A_{n}^{T})_{n\geq 1}$ is well-defined since $\left\lVert\eta^{T}\right\rVert_{1}<1$.
Let us introduce our main result. We consider for $t\in[0,1]$ the renormalized cumulative INAR($\infty$) process
$Y_{t}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\sum_{s=1}^{\lfloor Tt\rfloor}X_{s}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}N_{\lfloor Tt\rfloor}^{T},$ (5)
and its associated cumulative intensity
$\Lambda_{t}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\sum_{s=1}^{\lfloor Tt\rfloor}\lambda_{s}^{T}.$ (6)
The martingale $Z^{T}$ is defined as
$Z_{t}^{T}:=\sqrt{\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}}(Y_{t}^{T}-\Lambda_{t}^{T}).$ (7)
Recall the definition of the function $f^{\alpha,\lambda}$ from Section 3.1 in Jaisson and Rosenbaum (2016), which is
$f^{\alpha,\lambda}(x)=\lambda x^{\alpha-1}E_{\alpha,\alpha}(-\lambda x^{\alpha})$ (8)
and
$F^{\alpha,\lambda}(t)=\int_{0}^{t}f^{\alpha,\lambda}(x)dx,$ (9)
where $E_{\alpha,\beta}$ is the Mittag-Leffler function, defined as
@@
For a comprehensive review of its properties and applications, see, e.g., Haubold et al. (2011). We are now ready to state our results about the convergence in distribution of $(Z^{T},Y^{T})$ under the skorohod topology. We first introduce the following lemma, a discrete version of Lemma 4.3 in Jaisson and Rosenbaum (2016), which will be useful in the proof of equation (12) and Proposition 1.
Lemma 1.

### v7

@@ Abstract
Keywords: heavy–tailed unstable INAR($\infty$) processes, fractional CIR processes, rough volatility, option pricing.
MSC: 60G22, 60F05, 91G20
@@ 1 Introduction
Let $(\epsilon_{n})_{n\in\mathbb{Z}}$ be i.i.d. Pois($\nu$) random variables. Let $(\eta_{k})_{k\geq 1}$ be a sequence of positive numbers in $\ell^{1}$. For $n\in\mathbb{Z}$, $k\geq 1$, and $l\geq 1$, let $\xi_{l}^{(n,k)}$ be mutually independent random variables with a Pois($\eta_{k}$) distribution, which are also independent of the sequence $(\epsilon_{n})_{n\in\mathbb{Z}}$.
@@
To establish a more compact notation, we introduce the reproduction operator, denoted by $\star$. For a reproduction parameter $\iota\geq 0$, we define the offspring sequence $(\xi_{n}^{(\iota)})_{n\geq 1}$ as a sequence of i.i.d. random variables, where each element $\xi_{n}^{(\iota)}$ is an offspring variable following a Poisson distribution with mean $\iota$.
@@
${\iota\star Y}:=\sum_{n=1}^{Y}\xi_{n}^{(\iota)},$
provided that $Y$ is independent of the offspring sequence $(\xi_{n}^{(\iota)})$. Note that when $\xi_{n}^{(\iota)}\sim\mathrm{Bernoulli}(\alpha)$, the operator $\iota\star Y$ coincides with the classical binomial thinning operator $\iota\circ Y$ used in the INAR($p$) literature. Using this operator, the system of stochastic difference equations can be compactly rewritten as:
$\epsilon_{n}=X_{n}-\sum_{k=1}^{\infty}{\eta_{k}\star X_{n-k}}.$
Finally, we call $\nu$ the immigration parameter, $(\epsilon_{n})$ the immigration sequence, and $\eta_{k}$ the reproduction coefficient for each non-negative integer $k$.
@@
$N_{n}=\sum_{s=1}^{n}X_{s}.$ (1.1)
Hawkes processes, introduced by Hawkes (1971), are continuous-time self-exciting point processes widely used in various fields. Discrete-time analogs, such as cumulative INAR($\infty$) processes, provide similar modeling capabilities, particularly for count data observed at fixed time intervals. However, the scaling limits of heavy-tailed, nearly unstable (i.e., near-critical) cumulative INAR($\infty$) processes differ considerably from those of continuous-time counterparts, which warrants a dedicated investigation. Here “nearly unstable” means that the total reproduction mean is close to the critical value $1$ (see (1.3)). Under certain conditions, the Poisson autoregressive process can be viewed as an INAR($\infty$) process with Poisson offspring. For a comprehensive discussion on Poisson autoregressive models and their connections to INAR and Hawkes processes, refer to Fokianos (2024), Huang and Khabou (2023) and Kirchner (2016). From Proposition 3 in Kirchner (2016), we see that if an INAR($\infty$) process $(X_{n})_{n\geq 1}$ starts from time $1$, it can be specified recursively through its conditional mean (or intensity) $(\lambda_{n})_{n\geq 1}$ as follows:
$\lambda_{n}:=\nu+\sum_{s=1}^{n-1}\eta_{n-s}X_{s},\qquad n\geq 1,$ (1.2)
where $\nu>0$ is the immigration rate, and $(\eta_{n})_{n\geq 1}\in\ell^{1}$ represents the offspring distribution with $\eta_{n}\geq 0$ for all $n\in\mathbb{N}$. Let $\mathcal{F}_{n}:=\sigma(X_{1},\dots,X_{n})$ (with $\mathcal{F}_{0}$ the trivial $\sigma$-field); when we later consider the $T$-dependent sequence $(X_{n}^{T})$, we use the same notation $\mathcal{F}_{n}:=\sigma(X_{1}^{T},\dots,X_{n}^{T})$ for the filtration it generates. Then $\lambda_{1}=\nu$ and, conditionally on $\mathcal{F}_{n-1}$, the count $X_{n}$ is Poisson with parameter $\lambda_{n}$, i.e.,
$X_{n}\mid\mathcal{F}_{n-1}\sim\text{Pois}(\lambda_{n}).$
@@
under the Skorokhod topology, with $\sigma$ an explicit constant and $(W_{t})$ a Brownian motion. This result has been extended in Zhu (2013) to the case of nonlinear Hawkes processes. Regarding unstable Hawkes processes, the light-tailed case and the heavy-tailed case are respectively studied in Jaisson and Rosenbaum (2015) and Jaisson and Rosenbaum (2016). The light-tailed case corresponds to a CIR limit obtained under the crucial assumption
@@
a sequence of appropriately scaled random step functions formed from the process converges weakly towards a squared Bessel process. In the light-tailed (summable-kernel) setting, closely related results are available and suggest a squared-Bessel-type scaling limit for INAR($\infty$). Indeed, unstable limits are proved for finite-order discrete-time models (INAR/INAP($p$)) in Barczy et al. (2011), and the nearly-unstable light-tailed scaling limit is established for continuous-time Hawkes processes in Jaisson and Rosenbaum (2015). To the best of our knowledge, a dedicated proof for the INAR($\infty$) light-tailed case is not explicitly available in the literature. Consequently, scaling limits for nearly-unstable INAR($\infty$) models with heavy-tailed reproduction kernels appear to be missing from the literature, and this is the gap we aim to fill.
In this paper, we aim to provide a discrete-time counterpart to the continuous-time results found in Jaisson and Rosenbaum (2016), focusing on heavy-tailed nearly unstable INAR($\infty$) processes. Specifically, we consider a baseline reproduction kernel $(\eta_{n})_{n\geq 1}$ with heavy tail
@@
for some $\alpha\in\left(\frac{1}{2},1\right)$ and $K>0$, and we study a sequence of INAR($\infty$) processes $(X_{n}^{T})_{n\in\mathbb{N}}$ as $T\to\infty$. Throughout, “nearly unstable” means that the total reproduction mean approaches the critical value $1$, namely
$\|\eta^{T}\|_{1}=\sum_{k\geq 1}\eta_{k}^{T}\uparrow 1\quad\text{as }T\to\infty.$ (1.3)
In our parametrization, we take $\eta^{T}=a_{T}\eta$ with $a_{T}<1$ and $a_{T}\uparrow 1$, so that $\|\eta^{T}\|_{1}=a_{T}$. For each $T$, $(X_{n}^{T})_{n\geq 1}$ is an INAR($\infty$) process with immigration parameter $\mu^{T}$ and reproduction coefficients $(\eta_{k}^{T})_{k\geq 1}$. We denote by $\lambda_{n}^{T}:=\mathbb{E}[X_{n}^{T}\mid\mathcal{F}_{n-1}]$ its conditional mean (intensity), where here $\mathcal{F}_{n-1}=\sigma(X_{1}^{T},\dots,X_{n-1}^{T})$. The sequence of parameters $(\lambda_{n}^{T})$ is defined by
$\lambda_{n}^{T}=\mu^{T}+\sum_{s=1}^{n-1}\eta_{n-s}^{T}X_{s}^{T},\quad n\in\mathbb{N},$ (1.4)
where $\mu^{T}$ is a sequence of positive real numbers. As in Jaisson and Rosenbaum (2016), we establish specific assumptions on $\eta^{T}$ to facilitate our analysis. We prove tightness of the renormalized cumulative INAR()($\infty$) process $Y^{T}$ and the associated martingale $Z^{T}$ defined in (2.2) and (2.4). Moreover, we identify any subsequential limit $(Z,Y)$ of $(Z^{T},Y^{T})$ as a continuous martingale $Z$ whose sharp bracket satisfies $Z,Z]_{t}=Y_{t}$, where $Y$ is an integrated fractional Cox–Ingersoll–Ross (CIR) process.
@@
    A Computationally Superior Simulation Framework: Perhaps the most significant practical contribution is that INAR($\infty$) processes are substantially easier to simulate than their continuous-time, singular-kernel Hawkes counterparts. Simulating our model only requires iterative sampling from a Poisson distribution. Our scaling limit theorem thus serves as a crucial bridge: it rigorously validates that this simple discrete simulation converges to the complex continuous process of interest. This directly addresses the question of “why not just use a Hawkes process?”: The INAR($\infty$) process provides a computationally tractable pathway for its analysis and application.
@@
    New Applications in Financial Modeling: This framework opens the door to new applications not explored in the original Hawkes process literature. For instance, as demonstrated in the recent work of Wang et al. (2025), a bivariate version of this INAR($\infty$) framework can be shown to converge to the rough Heston model. This provides not only a novel microstructural foundation for rough volatility but also a simulation engine for pricing both European and path-dependent options, a task that remains challenging for other methods.
Our primary result, encapsulated in Theorem 2.13, demonstrates that under heavy-tailed and unstable conditions, the cumulative INAR($\infty$) process converges to a rough fractional diffusion process. This finding is achieved by leveraging discrete renewal theory and martingale central limit theorems.
@@ 2 Assumptions and Main Results
Roadmap of Section 2.
We collect here the definitions and statements that will be used throughout the rest of the paper. Assumption 2.1 specifies the heavy-tailed baseline kernel $(\eta_{n})$ and the near-critical scaling $a_{T}\uparrow 1$, while Assumption 2.3 fixes the joint scaling of $(a_{T},\mu^{T})$ that leads to a non-degenerate limit. We then introduce the rescaled cumulative process $Y^{T}$ in (2.2), the rescaled cumulative intensity $\Lambda^{T}$ in (2.3), and the rescaled martingale $Z^{T}$ in (2.4); these are the main objects in the tightness and limit identification arguments. Lemma 2.4 identifies the limiting Mittag–Leffler kernel via Laplace transforms (used to identify the drift term in the limit), whereas Lemma 2.6 provides $C$-tightness inputs (used to obtain tightness of $(Z^{T},Y^{T})$ in Proposition 2.10). The remaining lemmas in this section are auxiliary tools: Lemma 2.2 and Lemma 2.5 are Tauberian ingredients used in the proof of Lemma 2.4, and Lemma 2.7 provides a uniform pointwise bound on the resolvent coefficients, used only in the moment/tightness estimates. Proofs are deferred to Section A and B.
2.1 Heavy-tail and scaling assumptions
We normalize the baseline kernel so that $\|\eta\|_{1}=1$, and we propose two assumptions similar to those in Jaisson and Rosenbaum (2016).
Assumption 2.1.
@@ 2.1 Heavy-tail and scaling assumptions
    The baseline offspring distribution $(\eta_{n})_{n\geq 1}$ is a probability mass function on $\mathbb{N}$, i.e.
    $\eta_{n}\geq 0,\qquad\sum_{n=1}^{\infty}\eta_{n}=1,$
    and it has a heavy tail in the sense that
    $\eta_{n}\sim\frac{K}{n^{1+\alpha}}\quad\text{as }n\to\infty,$
    for some $\alpha\in\left(\frac{1}{2},1\right)$ and some constant $K>0$.
@@
Assumption 2.1(i) can equivalently be expressed in terms of the tail sum of $(\eta_{n})$. To this end, we state the following discrete Karamata-type lemma, which can be derived from the integral versions in (Bingham et al., 1989, Prop. 1.5.8 and Prop. 1.5.10) via a standard sum–integral comparison.
Lemma 2.2.
Let $l(t)$ be a positive slowly varying function that is locally bounded on $X_{0},\infty)$ for some $X_{0}>0$, and let $\beta\in\mathbb{R}$. For any fixed integer $X\geq X_{0}$, the following holds as integers $x\to\infty$:
  1. 1.
    If $\beta>-1$, then
    $\sum_{t=X}^{x}t^{\beta}l(t)\sim\frac{1}{\beta+1}x^{\beta+1}l(x).$
  2. 2.
    If $\beta<-1$, then
    $\sum_{t=x+1}^{\infty}t^{\beta}l(t)\sim-\frac{1}{\beta+1}x^{\beta+1}l(x).$}
Proof.
The proof is provided in Section A.1. ∎
Applying Lemma 2.2 with $\beta=-1-\alpha$ and $l(\cdot)\equiv K$ yields
$\sum_{k=n+1}^{\infty}\eta_{k}\sim\frac{K}{\alpha}\,n^{-\alpha}\quad\text{as}\quad n\to\infty,$}
and hence
@@
We now specify the asymptotic regime for $(a_{T})$ and $(\mu^{T})$. Define the constant
$\delta\;:=\;K\,\frac{\Gamma(1-\alpha)}{\alpha},$
which will naturally arise from the small-$z$ expansion of the Laplace transform of the heavy tail (see Lemma 2.4).
Recall from (1.4) that the (conditional) intensity sequence $(\lambda_{n}^{T})_{n\geq 1}$ is given by
$\lambda_{n}^{T}\;=\;\mu^{T}+\sum_{s=1}^{n-1}\eta_{n-s}^{T}X_{s}^{T},\qquad n\geq 1,$
where $\mu^{T}>0$ is the (possibly $T$-dependent) immigration parameter and $(\eta_{k}^{T})_{k\geq 1}\in\ell^{1}$ are reproduction coefficients with $\eta_{k}^{T}\geq 0$. We define the filtration $\mathcal{F}_{n}:=\sigma(X_{1}^{T},\dots,X_{n}^{T})$ (with $\mathcal{F}_{0}$ trivial). Then, by construction, $\lambda_{n}^{T}=\mathbb{E}[X_{n}^{T}\mid\mathcal{F}_{n-1}]$ and
$X_{n}^{T}\mid\mathcal{F}_{n-1}\sim\mathrm{Pois}(\lambda_{n}^{T}),\qquad n\geq 1,$
so $(X_{n}^{T}-\lambda_{n}^{T})_{n\geq 1}$ is a martingale difference sequence. We also define the counting (cumulative) process $N_{n}^{T}:=\sum_{s=1}^{n}X_{s}^{T}$.
Assumption 2.3.
@@
where $\delta=K\frac{\Gamma(1-\alpha)}{\alpha}$. The sequence $(\mu^{T})$ is introduced in (1.4).
Rescaled processes and martingales.
Define the compensated martingale
$M_{n}^{T}:=\sum_{s=1}^{n}\left(X_{s}^{T}-\lambda_{s}^{T}\right)=N_{n}^{T}-\sum_{s=1}^{n}\lambda_{s}^{T},\qquad n\geq 1.$
Indeed, since $\lambda_{s}^{T}=\mathbb{E}[X_{s}^{T}\mid\mathcal{F}_{s-1}]$, we have $\mathbb{E}[X_{s}^{T}-\lambda_{s}^{T}\mid\mathcal{F}_{s-1}]=0$, hence $(M_{n}^{T})_{n\geq 0}$ is an $(\mathcal{F}_{n})$-martingale.
We also introduce the renewal (resolvent) sequence $(A_{n}^{T})_{n\geq 1}\in\ell^{1}$:
$A_{n}^{T}:=\sum_{k=1}^{\infty}(\eta^{T*k})_{n},$ (2.1)
where * denotes the discrete convolution
@@
and $\eta^{*k+1}=\eta*\eta^{*k}$. This is well-defined since $\|\eta^{T}\|_{1}=a_{T}<1$.
We consider, for $t\in[0,1]$, the rescaled cumulative process
$Y_{t}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\sum_{s=1}^{\lfloor Tt\rfloor}X_{s}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}N_{\lfloor Tt\rfloor}^{T},$ (2.2)
and its rescaled cumulative intensity
$\Lambda_{t}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\sum_{s=1}^{\lfloor Tt\rfloor}\lambda_{s}^{T}.$ (2.3)
The associated (rescaled) martingale is
$Z_{t}^{T}:=\sqrt{\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}}(Y_{t}^{T}-\Lambda_{t}^{T}).$ (2.4)
To justify the terminology, let $\mathcal{F}_{n}:=\sigma(X_{1}^{T},\dots,X_{n}^{T})$. Since $\lambda_{n}^{T}=\mathbb{E}[X_{n}^{T}\mid\mathcal{F}_{n-1}]$, the sequence $(X_{n}^{T}-\lambda_{n}^{T})_{n\geq 1}$ is a martingale difference sequence, and hence
$M_{n}^{T}:=\sum_{s=1}^{n}(X_{s}^{T}-\lambda_{s}^{T})$
is an $(\mathcal{F}_{n})$-martingale. Therefore,
$Y_{t}^{T}-\Lambda_{t}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\,M_{\lfloor Tt\rfloor}^{T}$
is a martingale in $t$, and so $(Z_{t}^{T})_{t\in[0,1]}$ is a deterministically rescaled martingale.
Mittag–Leffler kernel.
Recall the function $f^{\alpha,\lambda}$ from Section 3.1 in Jaisson and Rosenbaum (2016), defined for $x>0$ by
$f^{\alpha,\lambda}(x)=\lambda x^{\alpha-1}E_{\alpha,\alpha}(-\lambda x^{\alpha}),$ (2.5)
and for $t\geq 0$ by
$F^{\alpha,\lambda}(t)=\int_{0}^{t}f^{\alpha,\lambda}(x)\,dx.$ (2.6)
It is known that $f^{\alpha,\lambda}$ is a probability density on $(0,\infty)$ and $F^{\alpha,\lambda}$ is the corresponding distribution function (the Mittag–Leffler distribution). Here $E_{\alpha,\beta}$ is the Mittag–Leffler function
@@
We will repeatedly use its Laplace transform: for $z\geq 0$,
$\int_{0}^{\infty}e^{-zx}f^{\alpha,\lambda}(x)\,dx=\frac{\lambda}{\lambda+z^{\alpha}},$ (2.7)
equivalently, $\int_{0}^{\infty}e^{-zx}\,dF^{\alpha,\lambda}(x)=\lambda/(\lambda+z^{\alpha})$.
2.2 A key discrete Laplace limit and tightness
We first state the discrete analog of Lemma 4.3 in Jaisson and Rosenbaum (2016).
Lemma 2.4.